OpenStudy (anonymous):

Write the expression as the sine, cosine, or tangent of an angle. sin 9x cos x - cos 9x sin x

OpenStudy (anonymous):

here are my choices sin 10x cos 8x sin 8x cos 10x

OpenStudy (tkhunny):

Have you considered the identity: \(\sin(a-b) = \sin(a)\cos(b) - \cos(a)\sin(b))\)

OpenStudy (anonymous):

no i really dont understand

OpenStudy (tkhunny):

I do not believe it. Why have you been given this problem if you have no way to solve it? Look at that formula I wrote, especially the right-hand side. Compare it to your problem statement. You must see the similarity or you cannot solve this problem.

OpenStudy (anonymous):

i really just need the answer im am being timed and my class is about to end please help

OpenStudy (tkhunny):

If you will just look at it, you will be done. Stop telling yourself you can't do it. As soon as you see the similarity, the solution will present itself.

OpenStudy (anonymous):

yea i found the answer no thanks to you

OpenStudy (tkhunny):

You mean someone did the work for you. Oh well, maybe you'll learn from that.